Show necessary step for the following questions
("For the given matrix"
A=the first row [1 -1] second row [3 2]
B=the first row [3 -2] second row [0 1]

1) show that (A^t)^t=A
2) show that (A+B)^t=A^t+B^t
3) show that (4A)^t=4(A^t)
4) show that (AB)^t=A^tB^t)
1
Expert's answer
2020-09-15T17:30:41-0400
"A=\begin{bmatrix}
 1 & -1 \\
 3 & 2
\end{bmatrix}, B=\begin{bmatrix}
 3 & -2 \\
 0 & 1
\end{bmatrix}"

1)


"A^T=\begin{bmatrix}
 1 & 3 \\
 -1 & 2
\end{bmatrix}"

"(A^T)^T=\begin{bmatrix}
 1 & -1 \\
 3 & 2
\end{bmatrix}=A"

2)


"A+B=\begin{bmatrix}
 1 & -1 \\
 3 & 2
\end{bmatrix}+\begin{bmatrix}
 3 & -2 \\
 0 & 1
\end{bmatrix}="

"=\begin{bmatrix}
 1+3 & -1-2 \\
 3+0 & 2+1
\end{bmatrix}=\begin{bmatrix}
 4 & -3 \\
 3 & 3
\end{bmatrix}"

"(A+B)^T=\begin{bmatrix}
 4 & 3 \\
 -3 & 3
\end{bmatrix}"

"A^T=\begin{bmatrix}
 1 & 3 \\
 -1 & 2
\end{bmatrix}, B^T=\begin{bmatrix}
 3 & 0 \\
 -2 & 1
\end{bmatrix}"

"A^T+B^T=\begin{bmatrix}
 1 &3 \\
 -1 & 2
\end{bmatrix}+\begin{bmatrix}
 3 & 0 \\
 -2 & 1
\end{bmatrix}="

"=\begin{bmatrix}
 1+3 & 3+0 \\
 -1-2 & 2+1
\end{bmatrix}=\begin{bmatrix}
 4 & 3\\
 -3 & 3
\end{bmatrix}=(A+B)^T"

3)


"4A=4\begin{bmatrix}
 1 & -1 \\
 3 & 2
\end{bmatrix}=\begin{bmatrix}
 4(1) & 4(-1) \\
 4(3) & 4(2) 
\end{bmatrix}=\begin{bmatrix}
 4 & -4 \\
 12 & 8
\end{bmatrix}"

"(4A)^T=\begin{bmatrix}
 4 & 12 \\
 -4 & 8
\end{bmatrix}"

"A^T=\begin{bmatrix}
 1 & 3 \\
 -1 & 2
\end{bmatrix}"

"4A^T=4\begin{bmatrix}
 1 & 3 \\
 -1 & 2
\end{bmatrix}=\begin{bmatrix}
 4(1) & 4(3) \\
 4(-1) & 4(2) 
\end{bmatrix}=\begin{bmatrix}
 4 & 12\\
 -4 & 8
\end{bmatrix}"

"(4A)^T=\begin{bmatrix}
 4 & 12 \\
 -4 & 8
\end{bmatrix}=4A^T"

4)


"AB=\begin{bmatrix}
 1 & -1 \\
 3 & 2
\end{bmatrix}\cdot\begin{bmatrix}
 3 & -2 \\
 0 & 1
\end{bmatrix}="

"=\begin{bmatrix}
 1(3)+(-1)(0) & 1(-2)+(-1)(1) \\
 3(3)+2(0) & 3(-2)+2(1)
\end{bmatrix}=\begin{bmatrix}
 3 & -3 \\
 9 & -4
\end{bmatrix}"

"(AB)^T=\begin{bmatrix}
 3 & 9 \\
 -3 & -4
\end{bmatrix}"

"B^TA^T=\begin{bmatrix}
 3 & 0 \\
 -2 & 1
\end{bmatrix}\cdot\begin{bmatrix}
 1 & 3 \\
 -1 & 2
\end{bmatrix}="

"=\begin{bmatrix}
 3(1)+0(-1) & 3(3)+0(2) \\
 -2(1)+1(-1) & -2(3)+1(2)
\end{bmatrix}=\begin{bmatrix}
 3 & 9 \\
 -3 & -4
\end{bmatrix}"

"(AB)^T=B^TA^T"